OpenStudy (anonymous):

1. Three sides of a pentagon have a length of 26 cm each. Each of the remaining two sides has a length of 14.5 cm. What is the perimeter of the pentagon? A. 75 cm B. 81 cm C. 107 cm D. 377 cm 2. What is the perimeter of a rectangle with a width of 23.6 cm and a length of 52.9 cm? A. 76.5 cm B. 100.1 cm C. 153 cm D. 180 cm 3. Which equation can be used to find the perimeter of a regular octagon with sides of length 12 m? A. P = 8 + 12 B. P = 8(12) C. P = 12 ÷ 8 D. P = 2(8) + 2(12)

OpenStudy (linda3):

1.) First we find that a pentagon has 5 sides. After that we understand that finding perimeter means adding all the 5 sides together. So here's our formula S1+S2+S3+S4+S5= P for perimeter It tells us that the first 3 sides are 26 cm. So we fill that in our formula 26+26+26+S4+S5= P If we continue reading we discover that the other to sides are of 14.5 cm. So we plot those numbers in and add all the numbers together to get P. 26+26+26+14.5+14.5 = 107

OpenStudy (linda3):

107 cm *

OpenStudy (linda3):

2.) Now the perimeter of a rectangle is the same concept as the pentagon by adding all the sides together. With a rectangle you have Length and With. You will often see the formula as , 2L + 2W = P because it is the exact same thing as L+L+W+W= P knowing this we can plot in our numbers with the numbers given in the question, L= 52.9 and W= 23.6 52.9+52.9+23.6+23.6 = P we add the all the numbers together and get 153 cm

OpenStudy (linda3):

3.) Now what we need to know is that an octagon has 8 sides. An easy way to remember this is that an octopus has an 8 legs if your one of those people who have a hard time memorizing. knowing that it has 8 sides and all the sides are a length of 12 m. Like in the previous question we saw that another way of writing the formula that has more than one of the same lengths, we could right the number of those sides and times it by the length. So the formula for this question would be 8 (because there is 8 sides) x 12(because that is the length of each side) So your results would be P= 8(12)

OpenStudy (anonymous):

Oh my gosh thank you so so SO much!!!!

OpenStudy (linda3):

your welcome! I hope it makes more sense to you now ;-)

OpenStudy (anonymous):

Oh yes it does!